OpenStudy (anonymous):

Blake is setting up his tent at a renaissance fair. If the tent is 8 feet tall and the tether can be staked no more than 2 feet from the tent, how long should the tether be?

OpenStudy (anonymous):

 Here we are dealing with a right triangle, and the sides of the triangle are the height of the tent, and the ground, and then the tether makes up the hypotenuse. Thus by the pythagorean theorem: \[8^2+2^2= 68\] \[\sqrt{68}\approx8.246\]feet

OpenStudy (anonymous):

thank you... i need help finding the geometric mean of [4\sqrt{3}\] and [10\sqrt{3}\]